# Dynamic similarity of vortex shedding in a superfluid flowing past a penetrable obstacle

### v1

@@ III.2 Effective Diameter
Figure 3: Time-averaged flow fields around a penetrable obstacle for $V_{0}/\mu=0.9$, $\sigma/\xi=20$, and $v_{0}/c_{\mathrm{s0}}=0.4$. (a) Condensate density $\overline{\rho}(\mathbf{r})$; the dashed circle indicates the obstacle’s $1/e^{2}$ radius. (b) Flow speed $|\overline{\bm{v}}(\mathbf{r})|$ and (c) irrotational flow speed $|\overline{\bm{v}_{\mathrm{irr}}}(\mathbf{r})|$. (d) Difference $|\overline{\bm{v}_{\mathrm{irr}}}(\mathbf{r})|-\overline{c_{\mathrm{s}}}(\mathbf{r})$, where $\overline{c_{\mathrm{s}}}(\mathbf{r})$ is the local speed of sound as determined by $\overline{\rho}(\mathbf{r})$. The black contour is the Mach-1 locus $|\overline{\bm{v}_{\mathrm{irr}}}(\mathbf{r})|=\overline{c_{\mathrm{s}}}(\mathbf{r})$.
Figure 4: Time-averaged vorticity fields $\overline{\omega_{\mathrm{s}}}(\mathbf{r})$ for various flow speeds $v_{0}/c_{\rm s0}$; (a-c) $V_{0}/\mu=0.9$ and $\sigma/\xi=20$, and (d-f) $V_{0}/\mu=0.9$ and $\sigma/\xi=40$. The black solid contour denotes the Mach-1 locus $\overline{v_{\rm irr}}(\mathbf{r})=\overline{c_{\rm s}}(\mathbf{r})$, enclosing the locally supersonic region, and the gray dotted circle marks the obstacle’s $1/e^{2}$ radius. The effective distance $D_{\mathrm{eff}}$ is defined as the maximum extent of the Mach-1 region along the $y$ direction. (g) Effective diameter $D_{\rm eff}/\xi$ versus vorticity width $w_{\rm v}/\xi$. The width $w_{\rm v}$ was determined from $\overline{\omega_{\mathrm{s}}}(\mathbf{r})$ (see text). Data points are shown for various values of $\sigma$ and $0.22\leq v_{0}/c_{\rm s0}\leq 0.65$ at $V_{0}/\mu=0.9$. Error bars are smaller than the symbols. The dashed line is a linear fit to the data.
Our analysis of the wake morphology and force response clearly shows a distinct dipole–to–cluster transition in the superfluid shedding dynamics for a penetrable obstacle. In this section, we introduce the effective diameter $D_{\rm eff}$ for a penetrable obstacle to determine the characteristic length scale relevant to the shedding dynamics, which would allow one to define a proper superfluid Reynolds number $\mathrm{Re}_{s}$ to organize the shedding behavior within the framework of dynamic similarity.
The nucleation of vortices in a superfluid occurs, in general, once compressibility effects become unstable [43, 44, 45, 46, 47]. This condition is referred to as the local Landau criterion and is expressed as
@@
where $c_{\rm s}(\mathbf{r})$ denotes the local speed of sound. This proposes that the spatial region around the obstacle where the flow becomes locally supersonic can be considered as the effective spatial extent of the obstacle in terms of vortex shedding. We thus define the effective obstacle diameter $D_{\rm eff}$ as the transverse width of the locally supersonic region (along the yy–direction).
To gain insight into the flow structure around a penetrable obstacle, we examine the spatial distributions averaged over time of the density $\overline{\rho}(\mathbf{r})$ and velocity $\overline{\bm{v}}(\mathbf{r})$. The exemplary data in the case of $v_{0}>v_{\mathrm{th}}$ are shown in Fig. 3. The density of condensate $\overline{\rho}$ exhibits a slight suppression compared to that of a stationary flow behind the obstacle due to the trail of the shed vortices downstream (Fig. 3(a)). The average flow speed $|\overline{\bm{v}}|$ in Fig. 3(b) clearly reflects the influence of these vortices: downstream of the obstacle, a low–velocity region develops where the background flow is partially canceled by the vortex-induced flow, while on the upstream side of the obstacle center the flow velocity is locally enhanced.
Given that we are interested in determining the supersonic region before vortex shedding, it is desirable to estimate the velocity field without the chaotic wake induced by shed vortices. This can be done by calculating the compressible (irrotational) component of the velocity field $\bm{v}_{\rm irr}(\mathbf{r})$ [23, 25, 44]. According to Helmholtz decomposition, $\bm{v}=\bm{v}_{\mathrm{irr}}+\bm{v}_{\mathrm{sol}}$ with $\nabla\times\bm{v}_{\mathrm{irr}}=0$ and $\nabla\cdot\bm{v}_{\mathrm{sol}}=0$. Because the contribution from vortices is isolated in the solenoidal component $\bm{v}_{\mathrm{sol}}$, $\bm{v}_{\rm irr}$ cleanly separates the compressibility-driven instability from vortex-induced fluctuations. Because vortex nucleation is triggered by compressibility-related instabilities, it is physically more relevant to base $D_{\rm eff}$ on the compressible component.
In Fig. 3(c), we observe that the time–averaged irrotational flow speed reaches its maximum at the obstacle center. This means that phase slips caused by vortex nucleation, which relax the accumulated phase gradient, lead to an overall reduction of $|\overline{\bm{v}(\rm r)}|$ as shown in Fig. 3(b). Consequently, we define the relevant supersonic region as
@@
with $\overline{c_{s}}(\mathbf{r})=\sqrt{g\overline{\rho}(\mathbf{r})/m}$ and determine $D_{\rm eff}$ to be the transverse extent of $\mathcal{S}$ through the obstacle center. In other words, $D_{\mathrm{eff}}$ corresponds to the transverse size of the Mach–1 contour around the obstacle (black solid line in Fig. 3(d)).
To test whether the supersonic region identified through $\overline{\bm{v}_{\rm irr}}$ suitably represents the vortex nucleation region, we compare it with the spatial distribution of the vortices. Figs. 4(a)4(f) show the vorticity field averaged over time, $\overline{\omega_{\mathrm{s}}}(\mathbf{r})$, together with the supersonic region defined by Eq. (6). Trails of positive (blue) and negative (red) vorticity appear just downstream of the obstacle. As the velocity increases, the vorticity-concentrated region expands and the vortex trail behind the obstacle grows into an enlarged wake along with the expansion the supersonic region. $\overline{\omega_{\mathrm{s}}}$ is predominantly concentrated along the lateral boundaries of the supersonic zone [63], in contrast to the fixed Gaussian width of the obstacle.
In Fig. 4(g), we plot the effective diameter $D_{\rm eff}$ for various Gaussian widths $\sigma$ and flow speed $v_{0}$ compared to the lateral width $w_{\mathrm{v}}$ of the distribution $\overline{\omega_{\mathrm{s}}}(\mathbf{r})$. Here, $w_{\mathrm{v}}$ is obtained by integrating $\overline{\omega_{\mathrm{s}}}(\mathbf{r})$ along the xx–direction within a finite region that encompasses the supersonic zone and extends downstream by $20\xi$ and by then determining the distance between the positive and negative peaks in the projected $\omega_{\mathrm{s}}$ profile. The effective diameter shows a linear relationship with $w_{\mathrm{v}}$, $D_{\rm eff}\approx 1.31w_{\mathrm{v}}$, for a wide range of obtacle conditions. This shows that the supersonic region is the relevant length scale of the penetrable obstacle for vortex shedding. An analytical description of the $v_{0}$-dependence of $D_{\mathrm{eff}}$ is presented in Appendix A.
From this observation and motivated by Reeve et al. [25], we propose a superfluid Reynolds number for penetrable obstacles, as follows:
@@
In computing $\textrm{Re}_{\rm s}$, we determined the critical velocity $v_{c}$, following Refs. [20, 42], by using the imaginary-time propagation method to check the existence of a stationary ground-state solution for $v_{0}<v_{c}$.
@@ III.3 Superfluid Reynolds number $\text{\rm Re}_{\rm s}$ and Strouhal number St
Figure 6: Strouhal number $\mathrm{St}$ versus $\mathrm{Re}_{\rm s}$ for $V_{0}/\mu=0.7$ (a) and $0.9$ (b). $\mathrm{St}$ was computed using $D_{\text{eff}}$ [Eq. (8)]. Vertical dashed lines denote the dipole–to–cluster transition ($\rm Re_{s}=2$). Dotted curves are fits to $\mathrm{St}(\mathrm{Re}_{\rm s})=\mathrm{St}_{0}\!\left[1-\frac{A}{\mathrm{Re}_{\rm s}+B}\right]$; fit parameters are listed in each panel. Error bars represent the uncertainty in $\mathrm{St}$ from the finite numerical resolution in determining $D_{\mathrm{eff}}$ and the FFT frequency resolution to extract $f_{\mathrm{s}}$. (c) Asymptotic Strouhal number $\mathrm{St}_{0}$ as a function of $V_{0}/\mu$. $\mathrm{St}_{0}$ was obtained by averaging $\mathrm{St}$ over $0.55\leq v_{0}/c_{\mathrm{s0}}\leq 0.65$ where $\mathrm{St}$ is saturated ($\sigma/\xi=20$), and its error bar indicates the standard error of the mean.
Figs 6(a) and 6(b) present our numerical results for $V_{0}/\mu=0.7$ and 0.9, respectively, on the plane of $\mathrm{Re}_{s}$ and $\mathrm{St}$. For a fixed obstacle strength $V_{0}/\mu$, data points of various obstacle sizes collapse onto a single curve. Moreover, we observe that $\mathrm{St}$ remains nearly constant for $\mathrm{Re_{s}}>2$, representing a close analogy to the classical case where $\mathrm{St}$ saturates in the turbulent vortex shedding regime [4]. The $\mathrm{Re}_{\rm s}$$\rm St$ trend is well described by the empirical fit [2, 25]
@@
To understand how $\mathrm{St_{0}}$ depends on the obstacle strength, we extend our analysis across a wide range of barrier heights, including the impenetrable regime where $D_{\textrm{eff}}$ remains well-defined. As shown in Fig. 6(c), $\mathrm{St_{0}}$ gradually decreases from about $0.30$ at $V_{0}/\mu=0.7$ to around $0.15$ at $V_{0}/\mu=1.5$. Beyond this point, deeper in the impenetrable regime, $\mathrm{St_{0}}$ remains nearly constant at $\sim 0.15$, consistent with previous findings where the geometric obstacle diameter was used [25].
The systematic increase of $\mathrm{St_{0}}$ with a decrease of $V_{0}/\mu$ in the penetrable regime likely reflects the decreasing robustness of vortex clustering for weaker obstacles. Because the value of $\mathrm{St}$ characterizes the geometric spacing of vortex clusters in the wake, larger values of $\mathrm{St}$ correspond to shorter inter-cluster distances ($\approx v_{0}/(2f_{s})$). This leads to stronger ovelaps between successive vortices with a finite core size, suppressing the formation of a stable vortex street. In our simulations, a well-defined shedding frequency ceases to appear for $V_{0}/\mu<0.6$, consistent with previous observations [41].
@@ III.4 Drag force
We note that this particular choice of drag coefficient is not unique [49].
@@
To quantify the dependence of $C_{\mathrm{D}}$ on $\mathrm{Re_{s}}$, we fit a power-law function of $C_{\mathrm{D}}=b\,\mathrm{Re_{s}}^{-\alpha}$ to the data for the two different ranges, $\mathrm{Re_{s}<2}$ and $\mathrm{Re_{s}}\geq 2$, separately (Fig. 7 insets). For the low-$\rm Re_{s}$ dipole shedding regime, the exponent obtained is in the range of $\alpha=0.6\sim 0.7$, similar with Stokes’ law in the 2D cylinder case in classical fluid dynamics [50, 51]. For the high-$\mathrm{Re}$ cluster shedding regime, the exponent decreases to $\alpha=0.22\sim 0.27$. Although $C_{\mathrm{D}}$ gradually decreases with an increase of $\mathrm{Re_{s}}$ in our parameter range, it is not known whether it vanishes as $\mathrm{Re_{s}}\rightarrow\infty$. In classical fluids, it has been argued that the dissipation rate remains finite as $\mathrm{Re\rightarrow\infty}$, implying a nonvanishing drag coefficient. However, studies under smooth wall or periodic boundary conditions have shown that the dissipation rate can decrease with an increase of $\mathrm{Re}$ [52, 53], suggesting that the drag coefficient may vanish. How this would be transformed in superfluid systems without intrinsic viscosity would be an intersesting problem.

### v2

@@ III.2 Effective Diameter
Figure 3: Time-averaged flow fields around a penetrable obstacle for $V_{0}/\mu=0.9$, $\sigma/\xi=20$, and $v_{0}/c_{\mathrm{s0}}=0.4$. (a) Condensate density $\overline{\rho}(\mathbf{r})$; the dashed circle indicates the obstacle’s $1/e^{2}$ radius. (b) Flow speed $|\overline{\bm{v}}(\mathbf{r})|$ and (c) irrotational flow speed $|\overline{\bm{v}_{\mathrm{irr}}}(\mathbf{r})|$. (d) Difference $|\overline{\bm{v}_{\mathrm{irr}}}(\mathbf{r})|-\overline{c_{\mathrm{s}}}(\mathbf{r})$, where $\overline{c_{\mathrm{s}}}(\mathbf{r})$ is the local speed of sound as determined by $\overline{\rho}(\mathbf{r})$. The black contour is the Mach-1 contour $|\overline{\bm{v}_{\mathrm{irr}}}(\mathbf{r})|=\overline{c_{\mathrm{s}}}(\mathbf{r})$.
Figure 4: Time-averaged vorticity fields $\overline{\omega_{\mathrm{s}}}(\mathbf{r})$ for various flow speeds $v_{0}/c_{\rm s0}$; (a-c) $V_{0}/\mu=0.9$ and $\sigma/\xi=20$, and (d-f) $V_{0}/\mu=0.9$ and $\sigma/\xi=40$. The black solid line denotes the Mach-1 contour $\overline{v_{\rm irr}}(\mathbf{r})=\overline{c_{\rm s}}(\mathbf{r})$, enclosing the locally supersonic region, and the gray dotted circle marks the obstacle’s $1/e^{2}$ radius. The effective diameter $D_{\mathrm{eff}}$ is defined as the maximum extent of the Mach-1 region along the $y$ direction. (g) Effective diameter $D_{\rm eff}/\xi$ versus vorticity width $w_{\rm v}/\xi$. The width $w_{\rm v}$ was determined from $\overline{\omega_{\mathrm{s}}}(\mathbf{r})$ (see text). Data points are shown for various values of $\sigma$ and $0.22\leq v_{0}/c_{\rm s0}\leq 0.65$ at $V_{0}/\mu=0.9$. Error bars are smaller than the symbols. The dashed line is a linear fit to the data.
Our analysis of the wake morphology and force response clearly shows a distinct dipole–to–cluster transition in the superfluid shedding dynamics for a penetrable obstacle. In this section, we introduce the effective diameter $D_{\rm eff}$ for a penetrable obstacle to determine the characteristic length scale relevant to the shedding dynamics, enabling the definition of a proper superfluid Reynolds number $\mathrm{Re}_{s}$ to organize the shedding behavior within the framework of dynamic similarity.
Vortex nucleation in a superfluid occurs when the local flow speed reaches or exceeds the local speed of sound [45, 41, 42, 43, 44]. This condition is referred to as the local Landau criterion and is expressed as
@@
where $c_{\rm s}(\mathbf{r})$ denotes the local speed of sound. This suggests that the spatial region around the obstacle where the flow becomes locally supersonic can be considered as the effective spatial extent of the obstacle in terms of vortex shedding. We thus define the effective obstacle diameter $D_{\rm eff}$ as the transverse width of the locally supersonic region (along the yy–direction).
To gain insight into the flow structure around a penetrable obstacle, we examine the spatial distributions averaged over time of the density $\overline{\rho}(\mathbf{r})$ and velocity $\overline{\bm{v}}(\mathbf{r})$. The exemplary data in the case of $v_{0}>v_{\mathrm{th}}$ are shown in Fig. 3. The density of condensate $\overline{\rho}$ exhibits a slight suppression compared to that of a stationary flow behind the obstacle due to the trail of the shed vortices downstream [Fig. 3(a)]. The average flow speed $|\overline{\bm{v}}|$ in Fig. 3(b) clearly reflects the influence of these vortices: downstream of the obstacle, a low–velocity region develops where the background flow is partially canceled by the vortex-induced flow, while on the upstream side of the obstacle center the flow velocity is locally enhanced.
Given that we are interested in determining the supersonic region before vortex shedding, it is desirable to estimate the velocity field without the chaotic wake induced by shed vortices. This can be done by calculating the compressible (irrotational) component of the velocity field $\bm{v}_{\rm irr}(\mathbf{r})$ [22, 24, 42]. According to Helmholtz decomposition, $\bm{v}=\bm{v}_{\mathrm{irr}}+\bm{v}_{\mathrm{sol}}$ with $\nabla\times\bm{v}_{\mathrm{irr}}=0$ and $\nabla\cdot\bm{v}_{\mathrm{sol}}=0$.Since the vortex contribution is confined to the solenoidal component $\bm{v}_{\mathrm{sol}}$, $\bm{v}_{\rm irr}$ provides a clear separation between compressibility-driven instabilities and vortex-induced fluctuations. As vortex nucleation is initiated by instabilities associated with compressibility, it is therefore more physically meaningful to define $D_{\rm eff}$ in terms of the compressible component.
In Fig. 3(c), we observe that the time–averaged irrotational flow speed reaches its maximum at the obstacle center. This means that phase slips caused by vortex nucleation, which relax the accumulated phase gradient, lead to an overall reduction of $|\overline{\bm{v}(\rm r)}|$ as shown in Fig. 3(b). Consequently, we define the supersonic region as
@@
with $\overline{c_{s}}(\mathbf{r})=\sqrt{g\overline{\rho}(\mathbf{r})/m}$ and determine $D_{\rm eff}$ to be the transverse extent of $\mathcal{S}$ through the obstacle center. In other words, $D_{\mathrm{eff}}$ corresponds to the transverse size of the Mach–1 contour around the obstacle [black solid line in Fig. 3(d)]. We note that there is an alternative way to determine an effective diameter $D_{\rm eff}$ from the supersonic region $\mathcal{S}$. For example, the size of $\mathcal{S}$ can be characterized in terms of its area $A_{\mathcal{S}}$, i.e., the equivalent diameter $2\sqrt{A_{\mathcal{S}}/\pi}$. A systematic comparison of such alternatives is left for future work.
To test whether the supersonic region identified through $\overline{\bm{v}_{\rm irr}}$ suitably represents the vortex nucleation region, we compare it with the spatial distribution of the vortices. Figures 4(a)4(f) show the vorticity field averaged over time, $\overline{\omega_{\mathrm{s}}}(\mathbf{r})$, together with the supersonic region defined by Eq. (6). Trails of positive (blue) and negative (red) vorticity appear just downstream of the obstacle. As the velocity increases, the vorticity-concentrated region expands and the vortex trail behind the obstacle grows into an enlarged wake along with the expansion of the supersonic region. $\overline{\omega_{\mathrm{s}}}$ is predominantly concentrated along the lateral boundaries of the supersonic zone [46], in contrast to the fixed Gaussian width of the obstacle.
In Fig. 4(g), we plot the effective diameter $D_{\rm eff}$ for various Gaussian widths $\sigma$ and flow speed $v_{0}$ compared to the lateral width $w_{\mathrm{v}}$ of the distribution $\overline{\omega_{\mathrm{s}}}(\mathbf{r})$. To obtain $w_{\mathrm{v}}$, we integrate $\overline{\omega_{\mathrm{s}}}(\mathbf{r})$ along the xx–direction within a finite region that encompasses the supersonic zone and extends downstream by $20\xi$. The width $w_{\mathrm{v}}$ is then defined as the distance between the positive and negative peaks in the projected $\omega_{\mathrm{s}}$ profile. The effective diameter shows a linear relationship with $w_{\mathrm{v}}$, $D_{\rm eff}\approx 1.31w_{\mathrm{v}}$, for a wide range of obstacle conditions, showing that the supersonic region sets the length scale of the penetrable obstacle for vortex shedding. An analytical description of the $v_{0}$-dependence of $D_{\mathrm{eff}}$ is presented in Appendix B.
From the observation of $D_{\rm eff}\propto w_{\mathrm{v}}$ and motivated by Reeves et al. [24], we propose a superfluid Reynolds number for penetrable obstacles, as follows:
@@
Physically, $\mathrm{Re_{s}}$ can be interpreted as the circulation of the background flow around the obstacle, measured in units of the circulation quantum $\kappa=h/m$, up to a numerical prefactor. This provides an estimate of the number of vortices that the flow can sustain around the obstacle. The replacement of $v_{0}$ by $v_{0}-v_{c}$ reflects the fact that dissipation is strictly absent below $v_{c}$, implicitly encoding obstacle-specific geometric details into the similarity parameter [24]. In computing $\textrm{Re}_{\rm s}$, we determined the critical velocity $v_{c}$, following Refs. [20, 34], by using the imaginary-time propagation method to check the existence of a stationary ground-state solution for $v_{0}<v_{c}$.
@@ III.3 Superfluid Reynolds number $\text{\rm Re}_{\rm s}$ and Strouhal number St
Figure 6: Strouhal number $\mathrm{St}$ versus $\mathrm{Re}_{\rm s}$ for $V_{0}/\mu=0.7$ (a) and $0.9$ (b). $\mathrm{St}$ was computed using $D_{\text{eff}}$ [Eq. (8)]. Vertical dashed lines indicate $\mathrm{Re}_{\mathrm{s}}=2$. Dotted curves are fits to $\mathrm{St}(\mathrm{Re}_{\rm s})=\mathrm{St}_{0}\!\left[1-\frac{A}{\mathrm{Re}_{\rm s}+B}\right]$; fit parameters are listed in each panel. Error bars represent the uncertainty in $\mathrm{St}$ from the finite numerical resolution in determining $D_{\mathrm{eff}}$ and the FFT frequency resolution to extract $f_{\mathrm{s}}$. (c) Asymptotic Strouhal number $\mathrm{St}_{0}$ as a function of $V_{0}/\mu$. $\mathrm{St}_{0}$ was obtained by averaging $\mathrm{St}$ over $0.55\leq v_{0}/c_{\mathrm{s0}}\leq 0.65$ where $\mathrm{St}$ is saturated ($\sigma/\xi=20$), and its error bar indicates the standard error of the mean.
Figures 6(a) and 6(b) present our numerical results for $V_{0}/\mu=0.7$ and 0.9, respectively, on the plane of $\mathrm{Re}_{s}$ and $\mathrm{St}$. For a fixed obstacle strength $V_{0}/\mu$, data points of various obstacle sizes collapse onto a single curve. Moreover, we observe that $\mathrm{St}$ remains nearly constant for $\mathrm{Re_{s}}>2$, representing a close analogy to the classical case where $\mathrm{St}$ saturates in the turbulent vortex shedding regime [4]. The $\mathrm{Re}_{\rm s}$$\rm St$ trend is well described by the empirical fit [2, 24]
@@
To understand how $\mathrm{St_{0}}$ depends on the obstacle strength, we extend our analysis across a wide range of barrier heights, including the impenetrable regime where $D_{\textrm{eff}}$ remains well-defined. As shown in Fig. 6(c), $\mathrm{St_{0}}$ gradually decreases from about $0.30$ at $V_{0}/\mu=0.7$ to around $0.15$ at $V_{0}/\mu=1.5$. Beyond this point, deeper in the impenetrable regime, $\mathrm{St_{0}}$ remains nearly constant at $\sim 0.15$, consistent with previous findings where the geometric obstacle diameter was used [24]. The smooth evolution of the asymptotic Strouhal number with $V_{0}/\mu$ suggests that the use of $\mathrm{Re}_{\rm s}$ provides a unified framework to characterize the penetrable-to-impenetrable crossover [20], which warrants future investigation.
The systematic increase of $\mathrm{St_{0}}$ with a decrease of $V_{0}/\mu$ in the penetrable regime likely reflects the decreasing robustness of vortex clustering for weaker obstacles. Because the value of $\mathrm{St}$ characterizes the geometric spacing of vortex clusters in the wake, larger values of $\mathrm{St}$ correspond to shorter inter-cluster distances ($\approx v_{0}/(2f_{s})$). This leads to stronger overlaps between successive vortices with a finite core size, suppressing the formation of a stable vortex street. In our simulations, a well-defined shedding frequency ceases to appear for $V_{0}/\mu<0.6$, consistent with previous observations [39]. For sufficiently weak obstacles, the flow disturbance is too small to form regular shedding and the critical velocity approaches $c_{s0}$ [28, 20].
@@ III.4 Drag force
We note that this particular choice of drag coefficient is not unique in the context of a superfluid [48].
@@
To quantify the dependence of $C_{\mathrm{D}}$ on $\mathrm{Re_{s}}$, we fit a power-law function of $C_{\mathrm{D}}=b\,\mathrm{Re_{s}}^{-\alpha}$ to the data for the two different ranges, $\mathrm{Re_{s}<2}$ and $\mathrm{Re_{s}}\geq 2$, separately [Fig. 7 insets]. For the low-$\rm Re_{s}$ dipole shedding regime, the exponent obtained is in the range of $\alpha=0.6\sim 0.7$, similar to Stokes’ law in the 2D cylinder case in classical fluid dynamics [50, 49]. For the high-$\mathrm{Re}$ cluster shedding regime, the exponent decreases to $\alpha=0.22\sim 0.27$. Although $C_{\mathrm{D}}$ gradually decreases with an increase of $\mathrm{Re_{s}}$ in our parameter range, it is not known whether it vanishes as $\mathrm{Re_{s}}\rightarrow\infty$. In classical fluids, it has been argued that the dissipation rate remains finite as $\mathrm{Re\rightarrow\infty}$, implying a nonvanishing drag coefficient. However, studies in smooth boundary or periodic boundary conditions have shown that the dimensionless dissipation can decrease with an increase of $\mathrm{Re}$ [51, 52], suggesting that the drag coefficient need not approach a finite constant at large $\mathrm{Re}$. How this behavior carries over to superfluids without intrinsic viscosity remains an interesting question.